# RealCam-I2V: Real-World Image-to-Video Generation with Interactive Complex Camera Control

### v1

@@ 4.1 Setup
Dataset. We train our model on RealEstate10K [78], which contains $\sim$ $70,000$ video clips with well-annotated camera poses. For metric depth alignment of absolute scene scale, we run COLMAP [40] point triangulator on each video clip with fixed camera intrinsics and extrinsics directly from RealEstate10K, obtaining the sparse point cloud of the reconstructed scene. We then calculate per-point depth scale against the metric depth from depth predictor. We term the median value of per-point depth scales in a frame as the frame-level depth scale. To make stable the training, we discard outliers of video clip whose maximum frame-level depth scale of the whole scene is among the top $2\%$ for too small values or the last $2\%$ for too large values, assuming sorted in ascending order. The same quantile filtering strategy is also applied on the minimum frame-level depth scales of video clips. It remains $58,000$ video clips for training and another $6,000$ for test. During training, we follow DynamiCrafter to sample $16$ frames from each single video clip while perform resizing, keeping the aspect ratio, and center cropping to fit in our training scheme. We train the model with a random frame stride ranging from $1$ to $10$ and take random condition frame as data augmentation. We fix the frame stride to $8$ and always use the first frame as the condition frame for inference.
Implementation Details. We choose DynamiCrafter [59] as our image-to-video (I2V) base model and seamlessly integrate proposed RealCam-I2V into it as a plugin. For metric depth predictor, we choose Depth Anything V2 [64] Large Indoor, which is fine-tuned on metric depth estimation. During depth-aligned training, we freeze all parameters of the base model and the depth predictor, while only parameters of proposed method are trainable. More details are listed in dataset section of Appendix. We supervise $\epsilon$-prediction on the model of $256\times 256$ resolution and $v$-prediction on the model of $512\times 320$ resolution respectively, following the pre-training scheme of DynamiCrafter. We apply the Adam optimizer with a constant learning rate of $1\times 10^{-4}$ with mixed-precision fp16 and DeepSpeed ZeRO-1. We train proposed method and variants on $8$ NVIDIA H100 GPUs with an effective batch size of $64$ for $50,000$ steps. More details are listed in implementation section of Appendix.
Rel. Abs. Rel. Abs. VideoGPT StyleGAN
@@
+ MotionCtrl [53] 1.0366 2.4932 5.8483 3.1244 6.2625 67.253 57.205
+ CameraCtrl [16]* 0.7373 1.7619 5.5090 2.1644 5.7648 69.202 58.900
+ CamI2V [77]* 0.4968 1.4853 3.4069 1.7253 3.5786 63.869 55.276
+ RealCam-I2V (Ours) 0.4052 1.3087 2.4709 1.4869 2.6095 55.229 48.080
w.r.t. CamI2V +18.44% +11.89% +27.47% +13.82% +27.08% +13.53% +13.02%
Table 1: Quantitative comparison with SOTA methods. * denotes the results we reproduced using DynamiCrafter as base I2V model. Our method achieves the state-of-the-art performance on both relative and absolute camera-controllable metrics, while coherently improve visual quality of generated videos, witnessed by a further drop of FVD. Best and second best results are highlighted respectively. We observe nearly $+30\%$ improvement on absolute metrics while over $+10\%$ improvement on relative metrics and FVD.
Method Quality Score I2V Score
Total subject background motion dynamic aesthetic imaging i2v i2v camera
Score consistency consistency smoothness degree quality quality subject background motion
Dynamicrafter (official) 85.25 94.74 98.29 97.83 40.57 58.71 62.28 97.05 97.56 20.92
Dynamicrafter (reproduced) 84.22 95.44 98.54 98.08 34.15 58.98 62.35 95.93 95.43 22.67
$+$ CamI2V 83.77 91.24 96.06 97.35 36.99 58.32 63.03 94.71 92.65 91.48
+++++ + Absolute Scale 84.37 90.99 96.23 97.36 46.75 58.37 62.91 94.73 93.44 87.42
+++++++ + + Noise Shaping (ours) 85.71 93.96 97.58 97.66 35.77 59.79 63.08 96.14 95.27 93.32
Table 2: Evaluation results on Vbench-I2V [23], a widely used benchmark suite with dynamic scenes and various types. Due to the efficient frozen parameters finetuning on Dynamicrafter, our method obtains the ability of camera control but decreases little in other metrics despite only training on static RealEstate10K.
@@ 4.2 Metrics
RotError. We calculate camera rotation errors by the relative angle between generated videos and ground truths in radians for rotation accuracy.
@@
TransError. For relative TransErr, we perform scene scale normalization on the camera positions of each video clip. The scene scale of generated video $\tilde{s}_{i}$ and grouth truth $s_{i}$ are individually calculated as the ${\cal L}_{2}$ distance from the first camera to the farthest one for each video clip. For absolute TransErr, we normalize both the video clip to the scene scale of ground truth video, i.e. $\tilde{s}_{i}=s_{i}$.
@@
CamMC. We perform the same scene scale normalization for relative metrics and absolute metrics as TransError, and evaluate the overall camera pose accuracy by directly calculating ${\cal L}_{2}$ similarity on camera-to-world matrices.
@@
FVD. We also assess the visual quality of generative videos by the distribution distance FVD [46] between generated videos and ground-truths.
@@ 4.3 Comparison with SOTA Methods
We compare our proposed method against models that either lack camera-condition training (DynamiCrafter [59]) or incorporate camera-condition training, namely DynamiCrafter+MotionCtrl [53] ($3\times 4$ camera extrinsics), DynamiCrafter+CameraCtrl [16] (Plücker embedding constructed from camera extrinsics and intrinsics as side input), and DynamiCrafter+CamI2V [77] (current SOTA using Plücker embedding and epipolar attention between all frames), as shown in Tab. 1.
Our method demonstrates significant improvements in visual quality (FVD) and camera control metrics (TransErr, RotErr, CamMC), particularly on absolute metrics. Specifically, absolute camera metrics improve by $+27\%$, relative camera metrics by $+14\%$, and FVD by $+13\%$. However, these improvements are not fully captured by the RealEstate10K dataset, which has limitations on movement speed and contains mostly static scenes. It’s strongly recommended to view the dynamic visualizations in the supplementary materials for a more comprehensive evaluation.
@@ 4.4 Ablation Study
Baseline RealCam-I2V Plugin RotErr $\downarrow$ TransErr $\downarrow$ CamMC $\downarrow$ FVD $\downarrow$
Absolute-Scale Scene-Constrained Noise Shaping Rel. Abs. Rel. Abs. VideoGPT StyleGAN
DynamiCrafter [53] 3.3415 9.8024 14.135 11.625 15.726 106.02 92.196
1.5163 6.6392 8.4607 7.2108 8.9505 71.942 65.014
+ MotionCtrl [53] 1.0527 2.2860 6.8182 2.9312 7.2272 70.292 60.845
0.8655 2.3342 4.2218 2.8083 4.5984 67.130 58.311
0.6373 2.0725 3.2308 2.3771 3.4721 58.885 50.111
+ CameraCtrl [16] 0.7373 1.7619 5.5090 2.1644 5.7648 69.202 58.900
0.7042 1.9477 3.8218 2.3007 4.0829 60.314 51.918
0.5436 1.7954 3.1845 2.0336 3.3620 55.004 46.702
+ CamI2V [77] 0.4968 1.4853 3.4069 1.7253 3.5786 63.869 55.276
0.4596 1.4109 3.0925 1.6282 3.2411 64.451 55.313
0.4052 1.3087 2.4709 1.4869 2.6095 55.229 48.080
Table 3: Ablation study. * denotes the results we reproduced using DynamiCrafter as base I2V model. Absolute scene-scale training resolves scale inconsistencies for real-world applications and its improvement on relative metrics indicates a more stable and unified camera control for video generation. Scene-constrained noise shaping can provides substantial improvements in dynamics and large camearabut is less effective than the combined approach, struggling with parameter tuning and dynamic consistency in lower noise stages. Best and second best results are highlighted respectively.
Figure 7: (a) Without Scene-Constrained Noise Shaping. (b) With Scene-Constrained Noise Shaping. The left shows failures in large movements without scene-constrained noise shaping, while the right illustrates a loss of dynamics when noise shaping is extended to lower noise stages.
Figure 8: Without kernel $\geq$3 in noise shaping, unvisible regions will be wrongly pasted to the generated video.
@@
Effect of absolute-scale training only. As shown in Tab. 3, compared to relative-scale training, absolute-scale training yields notable improvements, especially on absolute metrics. It implies that models trained on absolute-scale data can more accurately capture true-to-scale translations and better understand camera rotations within a realistic spatial framework. The absolute scene scale enhances robustness and compatibility, ensuring that the framework adapts effectively to real-world images and applications. This approach allows for interaction within a unified scale, enabling intuitive user control over camera actions.
@@
Effect of absolute-scale training + scene-constrained noise shaping. Adding scene-constrained noise shaping to a model trained with absolute-scale yields substantial gains in video quality and camera controllability. This improvement is evident across both camera metrics and FVD. The synergy of absolute-scale training and scene-constrained noise shaping ensures robust and precise control in diverse scenarios. As illustrated in Fig. 7, this combined approach delivers noticeably better dynamics compared to using scene-constrained noise shaping alone. Large camera movements, rotations, and rapid transitions, which previously struggled to maintain consistency and realism, now work seamlessly. This improvement underscores the strength of integrating absolute-scale training with noise shaping for complex motion scenarios.
@@
Effect of scene-constrained noise shaping only. As shown in Tab. 3, scene-constrained noise shaping can be used as the sole method for camera control when applied to a base model not trained with any camera conditions. It provides notable improvements in metrics, exemplified by nearly $50\%$ reduction on DynamiCrafter. However, this method underperforms compared to the combined method with absolute-scale training. It also introduces challenges in parameter selection. Applying shaping only in the high-noise phase limits camera contrl in lower noise stages, while extending shaping to mid-noise phase can suppress dynamic elements, resulting in static video output. This limitation affects the fluidity and responsiveness of generated camera movements, making the combination approach preferable for applications requiring natural dynamics.
@@ 4.5 Application
Figure 9: Visualization of various applications. Best viewed as dynamic videos in the supplementary materials.
As illustrated in Fig. 9, we demonstrate the versatility of our method through visualization results across various applications, including videos generated at resolutions of $512\times 320$ and $1024\times 576$ with camera control under complex scenarios, such as large movements or rotations. Additionally, our results include camera-controlled loop video generation, generative frame interpolation, and smooth scene transitions, highlighting the robustness of our approach. These visualizations showcase two major breakthroughs: first, our method achieves a real-world application breakthrough by addressing challenges like training-inference scale inconsistency and low usability, ensuring improved robustness and compatibility with real-world images. Second, our framework exhibits superior performance in complex camera motions, handling large and rapid movements, rotations, and dynamics more effectively than existing methods. More extensive results are provided in the supplementary materials.
@@ 5 Limitation Analysis and Future Work
The model was trained on datasets such as RealEstate10K, which consists primarily of real-world, indoor and outdoor videos collected from YouTube. This dataset’s content focuses heavily on realistic, static scenes, resulting in a model that excels in these contexts but performs less effectively when applied to scenes with significantly different visual styles, such as anime, oil paintings, or cartoon-like aesthetics. Better data quality, designing algorithm or improving the ability of fundamental model especially in long video generation will be considered into the future research.
6 Potential Negative Societal Impacts
The image-to-video generation technology developed in this work, with its enhanced camera controllability and breakthrough in real-world applications holds the potential for misuse, particularly in the creation of falsified or deceptive video content. The ability to precisely control camera movements and generate realistic sequences from single images could be exploited to produce convincing yet fabricated videos, leading to ethical concerns around misinformation and privacy violations. To mitigate these risks, we advocate for responsible usage and adherence to ethical guidelines when deploying the RealCam-I2V model.
@@ 7 Conclusion
Supplementary Material
We highly recommend that reviewers refer to index.html provided in our supplementary files.
@@ Appendix A Dataset
The camera trajectory of each video clip from RealEstate10K [^1] https://google.github.io/realestate10k/ [78] is first derived by SLAM methods at lower resolution with the field of view fixed at $90^{\circ}$. The authors then refine each of camera sequence using a structure-from-motion (SfM) pipeline, performing feature extraction, feature matching and global bundle adjustment successively. Given the unawareness of global scene scale, the resulted camera poses of RealEstate10K are up to an arbitrary scale per clip. For each frame the authors compute the $5$-th percentile depth among all point depths from that frame’s camera. Computing this depth across all cameras in a video clip gives a set of near plane depths and the whole scene is scaled so that the $10$-th percentile of this set of depths is $1.25$m.
@@
Appendix B Depth Predictor
We choose the metric depth version of Depth Anything V2222https://github.com/DepthAnything/Depth-Anything-V2/tree/main/metric_depth [64] as the metric depth predictor. Compared to their basic versions, the authors fine-tune the pre-trained encoder on synthetic datasets for indoor and outdoor metric depth estimation. The indoor model is capable of monocular metric depth estimation within a maximum depth of $20$m. We choose Large as the model size, which has $335.3$M parameters, and the indoor version. The scene scale of our model is aligned to the metric depth space of Depth Anything V2 Large Indoor, i.e. absolute scene scale.
Appendix C Training
We choose DynamiCrafter [^3] https://github.com/Doubiiu/DynamiCrafter [59] as our image-to-video (I2V) base model. We trained proposed method on $4$ publicly accessible variants of DynamiCrafter, namely 256, 512, 512_interp and 1024. We conduct ablation study on resolution $256\times 256$, due to the limitation of computing resource. For resolution $256\times 256$, we train all models on $\epsilon$-prediction with effective batch size $64$ on $8$ NVIDIA H100 GPUs for $50,000$ steps, taking about $25$ hours. For resolution $512\times 320$ and $1024\times 576$, we train RealCam-I2V on $v$-prediction while enable perframe_ae and gradient_checkpoint to reduce peak GPU memory consumption. We apply the Adam optimizer with a constant learning rate of $1\times 10^{-4}$ with mixed-precision fp16 and DeepSpeed ZeRO-1.

### v2

@@ 4.1 Setup
Dataset. We train our model on RealEstate10K [101], which contains 70,000 video clips with well-annotated relative-scale camera poses. For metric scene scale alignment, we run COLMAP [54] point triangulator on each video clip with fixed camera intrinsics and extrinsics directly from RealEstate10K, obtaining the sparse point cloud of the reconstructed scene. For metric depth predictor, we choose Depth Anything V2 [81] Large Indoor, which is fine-tuned on metric depth estimation. We then calculate per-point relative-to-metric scaling factor against metric depth prediction. We term the median value of scaling factors in a frame as the frame-level factor for robustness. To make stable training, we eliminate outliers of video clips whose maximum frame-level scaling factors are among the first 2% with too small values or the last 2% with too large values, assuming sorted in ascending order. The same quantile filtering strategy is applied on the minimum frame-level scaling factors of video clips. It remains 58,000 video clips for training and another 6,000 for testing. To further demonstrate the robustness of the proposed method while enhancing dynamics and diveristy, we reproduce RealCam-I2V on our open-sourced RealCam-Vid [98], a high-resolution video dataset with aligned metric-scale camera annotations and diverse scene dynamics. More experimental details can be found in  Appendix E.
Implementation Details. We choose DynamiCrafter [76] as our image-to-video (I2V) base model and seamlessly integrate proposed RealCam-I2V into it as a plugin. During depth-aligned training, we freeze all parameters of the base model and the depth predictor, while only parameters of proposed method are trainable. We follow DynamiCrafter to sample 16 frames from each single video clip while perform resizing, keeping the aspect ratio, and center cropping to fit in our training scheme. We train the model with a random frame stride ranging from 1 to 10 and take random condition frame as data augmentation. We fix the frame stride to 8 and always use the first frame as the condition frame for inference. We supervise $\epsilon$-prediction on the model of 256$\times$256 resolution and $v$-prediction on the model of 512$\times$320 resolution respectively, following the pre-training scheme of DynamiCrafter. We apply the Adam optimizer with a constant learning rate of $1\times 10^{-4}$ with mixed-precision fp16 and ZeRO-1.
Method MSA SNS RotErr $\downarrow$ TransErr $\downarrow$ CamMC $\downarrow$ FVD $\downarrow$
Relative Scale Metric Scale Relative Scale Metric Scale VideoGPT StyleGAN
@@
1.5163 6.6392 8.4607 7.2108 8.9505 71.942 65.014
MotionCtrl [69] 1.0527 2.2860 6.8182 2.9312 7.2272 70.292 60.845
0.8655 2.3342 4.2218 2.8083 4.5984 67.130 58.311
0.6373 2.0725 3.2308 2.3771 3.4721 58.885 50.111
CameraCtrl [22] 0.7373 1.7619 5.5090 2.1644 5.7648 69.202 58.900
0.7042 1.9477 3.8218 2.3007 4.0829 60.314 51.918
0.5436 1.7954 3.1845 2.0336 3.3620 55.004 46.702
CamI2V [97] 0.4120 1.3409 3.2934 1.5291 3.4230 62.439 53.361
0.4243 1.3596 2.6524 1.5539 2.8008 60.516 51.581
RealCam-I2V (Ours) 0.3884 1.2943 2.2317 1.4628 2.3609 53.718 45.460
Table 2: Ablation study of RealCam-I2V plugins. Metric Scene-scale Alignment (MSA) mitigates scale inconsistency for real-world applications, indicating a more stable and unified camera control. Scene-constrained Noise Shaping (SNS) solely provides substantial improvements on the base model but is less effective than the combined approach (ours). * denotes our reproduced results on DynamiCrafter [76]. Best and second best results are highlighted respectively.
Method Total I2V I2V Camera Subject Background Motion Dynamic Aesthetic Imaging
Score Subject Background Motion Consistency Consistency Smoothness Degree Quality Quality
  DynamiCrafter [76] 84.22 95.93 95.43 22.67 95.44 98.54 98.08 34.15 58.98 62.35
RealCam-I2V (Ours) 85.71 96.14 95.27 93.32 93.96 97.58 97.66 35.77 59.79 63.08
w.o. SNS 84.37 94.73 93.44 87.42 90.99 96.23 97.36 46.75 58.37 62.91
w.o. MSA 83.77 94.71 92.65 91.48 91.24 96.06 97.35 36.99 58.32 63.03
Table 3: Ablation study on Vbench-I2V [30], investigating how camera-conditioned fine-tuning exclusively on static RealEstate10K [101] data affects the generation quality and camera motion of the base model. Notably, even fine-tuned on a static dataset, it preserves dynamics (Dynamic Degree) without compromise. Introducing dynamic datasets will better enhance dynamics, we leave it for future work.
@@ 4.2 Metrics
RotError [22, 79, 97]. We calculate camera rotation errors by the relative angle between generated videos and ground truths in radians for rotation accuracy.
@@
TransError [22, 79, 97]. For relative TransErr, we perform scene scale normalization on the camera positions of each video clip. The scene scale of generated video $\tilde{s}_{i}$ and grouth truth $s_{i}$ are individually calculated as the ${\cal L}_{2}$ distance from the first camera to the farthest one for each video clip. For absolute TransErr, we normalize both the video clip to the scene scale of ground truth video, i.e. $\tilde{s}_{i}=s_{i}$.
@@
CamMC [69]. We perform the same scene scale normalization for relative and metric results as TransError, and evaluate the overall camera pose accuracy by directly calculating ${\cal L}_{2}$ similarity on camera-to-world matrices.
@@
FVD [61]. We also assess the visual quality of generative videos by the distribution distance between generated videos and ground truths.
@@ 4.3 Comparison with SOTA Methods
We compare RealCam-I2V with models that either lack camera-condition training (DynamiCrafter [76]) or incorporate camera-condition training, namely MotionCtrl [69] (3$\times$4 camera extrinsics), CameraCtrl [22] (Plücker embedding constructed from camera extrinsics and intrinsics as side input), and CamI2V [97] (current SOTA using Plücker embedding and epipolar attention between all frames), as shown in Tab. 1. We report all results using DynamiCrafter as the base model.
Our method demonstrates significant improvements in visual quality (FVD) and camera control metrics (TransErr, RotErr, CamMC), with particularly over 30% gains on metric scale results. However, these improvements are not fully captured by the RealEstate10K dataset, which contains mostly static scenes.
Figure 9: Visualization on various domains in real life scenarios. Despite training on RealEstate10K [101], our method can generalize naturally to out-of-domain images, including pets, landscape, anime, food and etc.
Figure 10: Visualization on complex trajectory, large movement and video dynamism. Our framework achieves precise trajectory adherence for complex camera motion paths while preserving high-fidelity video generation with dynamics.
@@ 4.4 Ablation Study
To validate our proposed method, we conducted a comprehensive ablation study on Metric Scene-scale Alignment (MSA) and Scene-constrained Noise Shaping (SNS).
@@
Effect of MSA only. As shown in Tab. 2, compared to relative scale training, metric scale training mitigates scale ambiguity in Fig. 4 and yields notable improvements, especially on metric scale results. It implies that models trained on metric-scale data can more accurately capture true-to-scale translations and better understand camera rotations within a realistic spatial framework. The metric scene scale enhances robustness and compatibility, ensuring that the framework adapts effectively to real-world input and applications. This approach allows for interaction within a unified scale, enabling intuitive user control over camera actions.
@@
Effect of MSA together with SNS. Adding scene-constrained noise shaping to a model with metric scale training yields substantial gains in video quality and camera controllability. This improvement is evident in both camera metrics and FVD. The synergy of metric-scale training and noise shaping ensures robust and precise control in diverse scenarios. This combined approach delivers noticeably better dynamics compared to using noise shaping alone. Large camera movements, rotations, and rapid transitions, which previously struggled to maintain consistency and realism, now work seamlessly. This improvement underscores the strength of integrating metric scale training with noise shaping for complex motion scenarios.
@@
Effect of SNS only. As shown in Tab. 2, scene-constrained noise shaping can be used as the only method for camera control when applied to a base model not trained with any camera conditions. It provides nearly 50% reduction on DynamiCrafter. However, this method performs poorly compared to the combined method with metric scale training. It also introduces challenges in parameter selection: Applying noise shaping only in the high noise stages limits camera control in lower noise stages, while extending noise shaping to mid-noise phase can suppress video dynamics, resulting in static video output. This limitation affects the fluidity and responsiveness of camera movements, making our combined approach preferable for applications requiring natural dynamics.
@@ 4.5 Application
As illustrated in Fig. 9 and Fig. 10, we demonstrate the versatility of our method through visualization results across various applications. Additionally, our results include camera-controlled loop video generation, generative frame interpolation, and smooth scene transitions, highlighting the robustness of our approach. These visualizations showcase two major breakthroughs: first, our method achieves a real-world application breakthrough by addressing challenges like training-inference scale inconsistency and low usability, ensuring improved robustness and compatibility with real-world images. Second, our framework exhibits superior performance in complex camera motions, handling large and rapid movements, rotations, and dynamics more effectively than existing methods.
@@ 5 Limitation Analysis and Future Work
Despite training in static RealEstate10K, as discussed in Fig. 9 and Fig. 10, our method demonstrate well generalization to videos with dynamic scenes, various styles and domains due to the designed algorithm to best preserve the knowledge of foundation model, However, the current model still suffers from real-world application mainly due to the limitation of data and there’s no suitable dataset designed for camera-controlled video generation, which requires both dynamic scenes and camera movement.
@@ 6 Conclusion
Appendix
Noise Shaping Subject Background Motion Dynamic Aesthetic Imaging I2V I2V Camera
Threshold Consistency Consistency Smoothness Degree Quality Quality Subject Background Motion
  / 90.99 96.23 97.36 46.75 58.37 62.91 94.73 93.44 85.85
900 93.96 97.58 97.66 35.77 59.79 63.08 96.14 95.27 93.32
800 95.02 97.91 97.79 26.42 60.07 63.33 96.53 95.73 95.15
600 95.10 98.02 97.80 27.64 60.21 63.85 96.55 95.72 96.72
Table 4: Evaluation results of noise shaping threshold $t_{\rm NS}$. For noise level $t\in[0,1000]$, the threshold $t_{\rm NS}$ denotes that we apply noise shaping only when $t>t_{\rm NS}$ in early denoising process.
Model Subject Background Motion Dynamic Aesthetic Imaging I2V I2V Camera
Consistency Consistency Smoothness Degree Quality Quality Subject Background Motion
  CogVideoX 1.5 [83] 91.80 94.66 97.07 40.98 62.29 70.21 96.46 95.50 39.71
w. RealCam-I2V (Ours) 97.81 98.41 99.33 44.31 64.31 70.70 99.03 99.40 91.35
Table 5: VBench-I2V results of RealCam-I2V trained on RealCam-Vid [98], with scene dynamics and large camera movements (near 360). The improvement in metrics can be attributed to the additional information from diverse camera traces in dynamic scenes. Noise shaping significantly improves the consistency and quality.
@@ Appendix A Dataset
The camera trajectory of each video clip from RealEstate10K [101] is first derived by SLAM methods at lower resolution with the field of view fixed at 90. The authors then refine each of camera sequence using a structure-from-motion (SfM) pipeline, performing feature extraction, feature matching and global bundle adjustment successively. Given the unawareness of global scene scale, the resulted camera poses of RealEstate10K are up to an arbitrary scale per clip. For each frame the authors compute the 5-th percentile depth among all point depths from that frame’s camera. Computing this depth across all cameras in a video clip gives a set of near plane depths and the whole scene is scaled so that the 10-th percentile of this set of depths is 1.25m.
@@
Appendix B Training
We choose DynamiCrafter [76] as our image-to-video (I2V) base model. We trained proposed method on $4$ publicly accessible variants of DynamiCrafter, namely 256, 512, 512_interp and 1024. We conduct ablation study on resolution 256$\times$256, due to the limitation of computing resource. For resolution 256$\times$256, we train all models on $\epsilon$-prediction with effective batch size 64 for 50,000 steps, taking about 50 hours. For resolution 512$\times$320 and 1024$\times$576, we train RealCam-I2V on $v$-prediction while enable perframe_ae and gradient_checkpoint to reduce peak GPU memory consumption. We apply the Adam optimizer with a constant learning rate of $1\times 10^{-4}$ with mixed-precision fp16 and ZeRO-1.